kprice:

At 1:00 P.M., you have 24 megabytes of a movie. At 1:15 P.M., you have 96 megabytes. What is the download rate in megabytes per minute? The download rate is megabytes per minute.

XxDeadxX:

do multiplacation

XxDeadxX:

not subtraction

kprice:

what do I multiply

HubbaBubba:

multiply 24 by 4 easy mafs

kprice:

tks

jhonyy9:

so from these given details result that 96-24 = 72 megabytes was downloaded in 15 minutes and now using the rule of 3 simple you can write 72 megabytes downloaded in 15 min. x MBytes in 1 min. so cross multiplie and calcule the x value x = 72*1/15 = 72/15 = ? MBytes hope helped understandably easy @supie